OpenStudy (anonymous):

The area of a triangle inscribed in a circle is 60m^2 and the radius of the circumscribed circle is 8cm. If the two sides of the inscribed triangle measure 10 cm and 12 cm, find the length of the third side.

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@jim_thompson5910 help me please :D

jimthompson5910 (jim_thompson5910):

I don't think it's possible, but I'm still trying various methods out.

OpenStudy (anonymous):

I think you can do herons formula here

jimthompson5910 (jim_thompson5910):

yeah you would solve the following for x \[\Large \sqrt{s(s-10)(s-8)(s-x)} = 60\] where \(\Large s = \frac{10+8+x}{2}\)

jimthompson5910 (jim_thompson5910):

the thing is though, I'm getting nonreal answers for x

jimthompson5910 (jim_thompson5910):

These are the approximate solutions the computer gives me -14.25982084+6.272359244i 14.25982084-6.272359244i -14.25982084-6.272359244i 14.25982084+6.272359244i none of which are a real number

jimthompson5910 (jim_thompson5910):

oh my bad, I typed it in wrong

jimthompson5910 (jim_thompson5910):

one second while I fix it

jimthompson5910 (jim_thompson5910):

ok, fixed solve the following for x \[\Large \sqrt{s(s-10)(s-12)(s-x)} = 60\] where \(\Large s = \frac{10+12+x}{2}\)

jimthompson5910 (jim_thompson5910):

The only issue I have with this now is that the triangle isn't 100% on the circle. I'm using geogebra and I can't get all 3 points to lie on the circle.

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

I worked on this problem last night. This thread shows what those before me had done. http://openstudy.com/users/directrix#/updates/54ecca39e4b06b22807cc9f1

OpenStudy (anonymous):

ok thnx

OpenStudy (anonymous):

can you help me with another problem

Directrix (directrix):

I am not sure that the current problem has been solved. The triangle that had to be right was off a bit.